Use the reduction of order method to solve the following problem given one of the solution y1.

(a) (x^2 - 1)y'' -2xy' +2y = 0 ,y1=x

(b) (2x+1)y''-4(x+1)y'+4y=0 ,y1=e^2x

(c) (x^2-2x+2)y'' - x^2 y'+x^2 y =0, y1=x

(d) Prove that if 1+p+q=0 than y=e^x is a solution of y''+p(x)y'+q(x)y=0, use this fact to solve (x-1)y'' - xy' +y =0
Answer #1

Solution: Let the second solution be y,(x) -ix Substitute these in (x2-1),--2xy, + 2y = O 1,(x,-x) = 21,, dr dr ー=2で-x) dz

dh Resubstitutez =-, dh Let ec = 1 and k=0 Let the second solution be y, (r)- ve 2 x + 1) y-4(x + 1) y, + 4y = 0. Substitut

4x: + (2x+1)-=0 dr :=-4 (2x +1)at dr In : =-2x-1+Inl2x+11-4C = A(2x+1)e-2x-1 dh Resubstitute - A(2x +1)e-1 v(x)=A(-e-a-h-e-2